# MAB-Based Channel Scheduling for Asynchronous Federated Learning in Non-Stationary Environments

### v1

@@ Abstract
Federated learning enables distributed model training across clients under central coordination without raw data exchange. However, in wireless implementations, frequent parameter updates between the server and clients create significant communication overhead. While existing research assumes known channel state information (CSI) or stationary distributions, practical wireless channels exhibit non-stationary characteristics due to channel fading, user mobility in telecommunication networks and hostile attacks. The unavailability of both CSI and time-varying statistics can lead to unpredictable failures in parameter transmission, exacerbating clients staleness thus affecting model convergence. To address these challenges, we propose an asynchronous federated learning scheduling framework for non-stationary channel environments, designed to reduce clients staleness while promoting both fair and efficient communication and aggregation. This framework considers two distinct channel scenarios: extremely non-stationary and piece-wise stationary channels. Age of Information (AoI) serves as a metric to quantify client staleness under non-stationary conditions. Firstly, we perform a rigorous convergence analysis to explore the impact of AoI and per-round client participation on learning performance. The channel scheduling problem under the non-stationary scenario is modeled as a multi-armed bandit (MAB) framework and we derive the achievable theoretical lower bounds on the AoI regret. Based on this framework, our research proposes corresponding scheduling strategies for the two non-stationary channel scenarios that leverage the foundations of the GLR-CUCB and M-exp3 algorithms, along with derivations of their respective upper bounds on AoI regret. Additionally, to address the issue of imbalanced client updates in non-stationary channels, we introduce an adaptive allocation strategy that incorporates considerations of marginal utility and fairness of clients. Simulation results demonstrate that the proposed algorithm achieves sub-linear growth in AoI regret, accelerates federated learning convergence, and promotes fairer aggregation.
@@ I Introduction
The proliferation of Internet of Things (IoT) devices and the rise of edge computing have resulted in an increasingly decentralized distribution of data across end devices, such as smartphones and sensors. In traditional centralized machine learning approaches, data consolidation at a single location is required, which raises privacy concerns and incurs significant communication overhead. In contrast, Federated Learning (FL) [1] offers a promising solution by enabling local training at the client side, where only model parameters or gradients are transmitted to a central server for aggregation, thereby preserving data privacy and reducing communication costs.
@@
Traditional synchronous federated learning (Sync-FL) faces inherent limitations, as the central server must await parameter updates from all selected clients before initiating model aggregation. This synchronous approach introduces a straggler effect [2] due to heterogeneous computing capabilities, unreliable network connections, and dynamic environments. To overcome these challenges, asynchronous federated learning (Async-FL) [3] has emerged, enabling the central server to perform aggregation upon receiving updates from any client, thus facilitating flexible client participation in the training process.
However, Async-FL introduces significant challenges in managing stale gradients and model performance. In Async-FL, clients with superior computational capabilities and communication conditions are able to participate in the aggregation of the global model frequently, while straggled clients participate sporadically. This update imbalance leads to model drift [4], where the global model diverges from the optimum and exhibits bias toward frequently participating clients. This drift becomes particularly significant with non-independently and identically distributed (non-i.i.d.) clients data [5, 6], while stale gradients information from straggled clients further impedes global model convergence.
Research by [7] analyzes the negative impact of stale gradients on convergence speed, while [8] extends this work by providing theoretical foundations for understanding how aged client gradients specifically affect FL systems. These findings highlight the importance of addressing staleness issues in Async-FL. Zhou et al. [9] proposed selecting stale gradients with a consistent descent direction to accelerate the overall training process. A semi-asynchronous federated aggregation (SAFA) approach[10] is proposed to aggregates and distributes the global model while allowing a certain latency. Yang et al. [11] introduced age of information(AoI) to quantify the staleness of client updates and proposed a scheduling policy that accounts for both AoI and instantaneous channel quality. Client scheduling strategies, which incorporate key metrics such as CSI, local data distribution, and AoI, have been extensively explored. Liu et al. [12] developed a resource allocation framework that integrates CSI and client data heterogeneity. Ozfatura et al. [13] leveraged average AoI across clients as a regularization term to address exclusive client scheduling in channel-aware federated learning. AoI-constrained optimization problems inherently constitute sequential decision-making scenarios, typically formulated as constrained Markov decision processes (CMDP)[5].
@@
While the aforementioned studies assume the availability of CSI and successful reception of model parameters from clients at the central server, practical scenarios involving IoT devices operating in unlicensed frequency spectrum bands (FSB) may face limitations in signal transmission capacity [14]. Furthermore, acquiring CSI via pilot transmission introduces additional signaling overhead, which increases as the number of clients grows [15]. Relevant studies typically operate under the stationary channel assumption, wherein the channel state is characterized by a steady but unknown distribution. Several studies have examined Federated Learning (FL) under unknown CSI. Amiri et al. [16] proposed receiver beamforming, while Razavikia et al. [17] used q-ary modulation without clients’ CSI. Tegin et al. [18] studied federated learning for over-the-air computing and utilized multiple antennas to mitigate the effects of time-varying fading channels. To address the challenge of unknown channels, researchers model multi-channel scheduling as a multi-armed bandit (MAB) problem, where users select channels with unknown distributions and receive rewards based on transmission success. Notably, existing studies have linked the cumulative reward to the freshness of the transmitted information. Bhandari et al. [19] introduced the problem of minimizing the cumulative AoI in a single-source, stationary independent and identically distributed (i.i.d.) channel system, framing it as AoI bandits with AoI regret as the optimization metric. Qian et al. [20] established policy-independent lower bounds on the average AoI for multiple users and channels. These studies assume that the channel is smooth, meaning the channel state follows a stationary but unknown distribution, or that the channel undergoes smooth fading.
@@
However, practical wireless environments exhibit inherently time-varying channel conditions [21], manifested through phenomena such as path fading in 5G millimeter-wave systems, user mobility in telecommunication networks [22], and hostile jamming in extremely non-stationary communication scenarios [23]. In federated learning systems, these time-varying channels can lead to unpredictable clients participation failures during global aggregation due to transmission delays exceeding prescribed communication deadlines. In non-stationary channels, the statistical properties may change abruptly due to network attacks or severe channel fading in mobile scenarios [24]. Mandal et al. [23] studied the multi-user scheduling problem in extremely non-stationary environments. For piecewise-stationary channels, the Combinatorial Upper Confidence Bound(CUCB) algorithm [25] was proposed to solve the combinatorial multi-armed bandit problem by balancing exploration and exploitation of unknown channels. Zhou et al. [26] proposed a change-point detector for detecting the moment when the channel undergoes a sudden change.
@@
Non-stationary channels can induce participation imbalance in aggregation, challenging federated learning fairness. This phenomenon is particularly prevalent in non-i.i.d data scenarios, where the global model exhibits bias towards frequently participating clients with superior channel conditions and transmission success rates. Zhu et al[27] proposed variance reduction to correct the inequality of participation rates of different clients. Federated learning fairness encompasses collaborative and performance dimensions. Collaborative fairness addresses contribution evaluation and reward allocation, considering non-iid data distribution’s impact on model convergence. Performance fairness targets model prediction bias reduction for enhanced cross-client generalization. In [28], the authors introduce a long-term fairness constraint for client selection to ensure that the average participation rate of each client does not fall below the expected guaranteed rate. CFFL[29] evaluates fairness by evaluating the validation performance of each client to represent reputation metrics.RRFL[30] considers the difference in the updates between each local model and the shared global model to represent the reputation value. CGSV [31] introduces the cosine gradient Shapley value to approximate the difference between client and global updates as the marginal utility. FedCE[32] estimates the client’s contribution in terms of gradient and data space. MAB-based client selection is used in [33] to balance latency and performance fairness.
@@
To the best of our knowledge, no research has explored the staleness and fairness problem of federated learning under the conditions where CSI is unavailable and the channels are non-stationary. We presents a MAB-based channel scheduling framework for federated learning in non-stationary wireless environments to optimize clients participation and accelerate global model convergence. The main contributions are summarized as follows:
@@
  • For asynchronous federated learning in non-stationary scenarios, we first conduct theoretical analysis, utilizing AoI to quantify client staleness. Our work provides quantitative analysis on the impact of clients’ AoI and successful participation rates and unifies these two factors to derive our optimization objective for channel scheduling. The channel scheduling problem under non-stationary channel conditions is modeled as an MAB framework. Our analysis encompasses two non-stationary channel scenarios: the piece-wise stationary case and the extremely non-stationary case.
@@
  • For both channel scenarios, we derive the theoretical lower bounds of AoI regret and propose corresponding channel scheduling strategies. For piecewise-stationary channels, we extend the GLR-CUCB algorithm, and for extremely non-stationary channels, we introduce the M-exp3 algorithm. For both strategies, we establish the asymptotic upper bounds of AoI regret. Our approaches achieve sublinear regret growth, effectively reducing client staleness in asynchronous federated learning.
@@
The organization of this paper proceeds as follows: Section II introduces the asynchronous federated learning model under non-stationary channels. Section III presents the convergence analysis. Section IV establishes the AoI regret lower bounds for both extremely non-stationary and piecewise-stationary channels, and derives the upper bound under the proposed strategy. Section V details the adaptive channel matching algorithm. Section VI evaluates the experimental results.
@@ II SYSTEM MODEL
This section presents the federated learning procedure and wireless channel model, focusing on two non-stationary channel types: extremely non-stationary and piece-wise stationary channels.
@@ II-A Procedure of Federated learning
We consider a non-stationary wireless network consisting of $M$ clients, indexed by the set =$\mathcal{M}=\{1,2,\dots,M\}$, and a central server. Each client $i\in\mathcal{M}$ holds a local dataset $\mathcal{D}_{i}$, with $|\mathcal{D}_{i}|$ representing the number of data samples. The clients collaborate with the central to complete privacy-preserving decentralized training, utilizing the collective dataset $D=\sum_{i=1}^{M}|\mathcal{D}_{i}|$. We denote the sample-wise loss function for client $i$ as $\mathcal{F}_{i}(\boldsymbol{w}_{i},x)$, where $\boldsymbol{w}_{i}$ represents the local model parameter of client $i$, and $x\in\mathcal{D}_{i}$ is a data sample from the dataset $\mathcal{D}_{i}$. Then the local loss function for client $i$ is given by:
@@
Due to the unknown and non-stationary nature of the channel CSI, only a subset of clients may successfully transmit their gradient updates to the central server in each round. We define $S_{t}$ as the set of clients that successfully transmit their updates in round $t$. Furthermore, we introduce the AoI to quantify the degree to which clients are out-of-sync in the global aggregation process. Specifically, Let $a_{i}(t)$ denote the AoI of client $i$ in aggregation round $t$ and $h_{i}(t)$ denotes the most recent round in which client $i$ successfully participated in the global aggregation before round $t$. We have
@@
$\bullet$ (Step 1 Global Model Broadcast) To mitigate computational resource wastage and optimize the utilization of local training outcomes, clients that failed to transmit their updates in round $t-1$ retain their gradient updates and do not participate in the new updates in round $t$. Therefore, in round $t$, the server only sends the latest global model parameters $\boldsymbol{w}_{t}$ to the clients in $S_{t-1}$. Considering the sufficient transmission capacity of the server, we assume that the downlink channel is always successful.
@@
The central server then assigns each client a channel to transmit its model update $\tilde{\boldsymbol{G}}_{i,t}$ to the server.
$\bullet$ (Step 4 Global Model Update) Due to the uncertainty in CSI and the non-stationary characteristics of the channels, only the model updates from clients with favorable channel conditions are successfully transmitted to the server. The global model can thus be updated as follows:
@@
where $\zeta_{i}^{t}$ denotes the aggregation weight of client $i$ in round $t$. The AoI of each client $i,\forall i\in\mathcal{M}$ is updated as
@@
The above steps are repeated until round $T$ is reached.
@@ II-B Non-stationary Wireless network
In this wireless network, we equally divide the spectrum into $N$ orthogonal sub-channels, indexed by the set $\mathcal{N}=\{1,2,\dots,N\}$. We consider a network with $M$ clients ($N\geq M$), and a central server. Each client periodically attempts to transmit the time-sensitive local model to the central server during global aggregation over one of the $N$ channels.
The central server allocates distinct sub-channels to clients based on a scheduling policy, preventing collisions within each round. Sub-channels are modeled as Bernoulli channels with state Good (1) or Bad (0) at any round $t$. Clients lack real-time CSI and prior knowledge of the channels’ statistical properties.
In the stationary setting, channel states are independent across channels over $T$ rounds. Let $\mu_{k}$ denote the mean state of channel $k\in\{1,2,\cdots,N\}$, which is unknown to client $i\in\{1,2,\cdots,M\}$ and constant throughout $T$. Without loss of generality, assume $\mu_{1}>\mu_{2}>\cdots>\mu_{N}$, where $\{1,\cdots,M\}$ are the $M$-best channels and $\{M+1,\cdots,N\}$ are the $M$-worst channels.
In the non-stationary setting, channel states may fluctuate due to factors like malicious attacks or collisions [23]. In the case of extreme non-stationarity, an adversary pre-determines the state sequence of each channel (Good or Bad) without relying on any statistical assumptions. In contrast, for piecewise-stationary channels, we assume the central server has some prior knowledge, where the mean channel state $\mu_{k}$ for each $k=1,2,\dots,N$ remains constant over certain intervals and changes only at unknown rounds.

### v2

@@ Abstract
Federated learning enables distributed model training across clients under central coordination without raw data exchange. However, in wireless implementations, frequent parameter updates between the server and clients create significant communication overhead. While existing research assumes either known channel state information (CSI) or that the channel follows a stationary distribution, practical wireless channels exhibit non-stationary characteristics due to channel fading, user mobility, and hostile attacks in telecommunication networks. The unavailability of both CSI and time-varying channel distribution can lead to unpredictable failures in parameter transmission, exacerbating clients staleness thus affecting model convergence. To address these challenges, we propose an asynchronous federated learning scheduling framework for non-stationary channel environments, designed to reduce clients staleness while promoting both fair and efficient communication and aggregation. This framework considers two channel scenarios: extremely non-stationary and piecewise-stationary channels. Age of Information (AoI) serves as a metric to quantify client staleness under non-stationary conditions. Firstly, we perform a rigorous convergence analysis to explore the impact of AoI and per-round client participation on learning performance. The channel scheduling problem in the non-stationary scenario is addressed and formulated within the multi-armed bandit (MAB) framework and we derive the achievable theoretical lower bounds on the AoI regret. Based on this framework, we propose corresponding scheduling strategies for the two non-stationary channel scenarios that leverage the foundations of the GLR-CUCB and M-exp3 algorithms, along with derivations of their respective upper bounds on AoI regret. Additionally, to address the issue of imbalanced client updates in non-stationary channels, we introduce an adaptive matching strategy that incorporates considerations of marginal utility and fairness of clients. Simulation results demonstrate that the proposed algorithm achieves sub-linear growth in AoI regret, accelerates federated learning convergence, and promotes fairer aggregation.
@@ I Introduction
The proliferation of Internet of Things (IoT) devices and the rise of edge computing have resulted in an increasingly decentralized distribution of data across end devices, such as smartphones and sensors. In traditional centralized machine learning approaches, data consolidation at a single location is required, which raises privacy concerns and incurs significant communication overhead. In contrast, federated learning (FL) [1] offers a promising solution by enabling local training on the client side, where only model parameters or gradients are transmitted to a central server for aggregation, preserving data privacy and reducing communication costs.
@@
Traditional synchronous federated learning (Sync-FL) faces inherent limitations, as the central server must await parameter updates from all selected clients before initiating model aggregation. This synchronous approach introduces the straggler effect [2] due to heterogeneous computing capabilities, unreliable network connections, and dynamic environments. To address these challenges, asynchronous federated learning (Async-FL) [3] has been introduced, where the central server aggregates as long as the model updates are received within a predefined time threshold in each round, enabling more flexible client participation.
However, Async-FL introduces significant challenges in managing stale gradients and model performance. In federated learning systems, the fundamental issue of client drift [4] arises from non-independently and identically distributed (non-i.i.d.) data across clients [5][6], where the global model can diverge from the optimum. This drift is intensified in Async-FL settings. Clients with superior computational capabilities and network conditions participate frequently in global model aggregation, whereas straggled clients contribute only sporadically. The resulting update imbalance amplifies the existing drift problem, creating a bias toward frequently participating clients. Furthermore, stale gradient information from straggled clients further impedes global model convergence. Dai et al. [7] analyzed the negative impact of stale gradients on convergence speed and suggests that the acceptable level of staleness in distributed training depends on model complexity. Zhou et al. [8] proposed selecting stale gradients with a consistent descent direction to accelerate the overall training process. A semi-asynchronous federated aggregation (SAFA) approach [9] was proposed to aggregate and distribute the global model while allowing a certain latency. Yang et al. [10] introduced age of information (AoI) to quantify the staleness of client updates and proposed a scheduling policy that accounted for both AoI and instantaneous channel quality. Ozfatura et al. [11] leveraged the average AoI between clients as a regularization term to address exclusive client scheduling in channel-aware federated learning.
@@
While the aforementioned studies assumed the availability of CSI and the successful reception of model parameters from clients on the central server, practical scenarios involving IoT devices operating in unlicensed frequency spectrum bands (FSB) face limitations in signal transmission capacity [12]. Furthermore, acquiring CSI via pilot transmission introduces additional signaling overhead, which increases as the number of clients grows [13]. Several studies examined federated learning under unknown CSI. Amiri et al. [14] considered the lack of CSI and designed a receiver beamforming scheme to compensate for it. In the context of federated edge learning, Razavikia et al. [15] employed q-ary quadrature amplitude modulation when the client does not have access to CSI. Tegin et al. [16] studied federated learning for over-the-air computing and utilized multiple antennas to mitigate the effects of time-varying fading channels. To address the challenge of unknown channel distributions, some studies modeled multi-channel scheduling as a multi-armed bandit (MAB) problem, where users schedule channels according to a strategy and receive rewards based on transmission success. Notably, existing studies linked the cumulative reward to the freshness of the transmitted information. Bhandari et al. [17] introduced the problem of minimizing the cumulative AoI in a single-source, stationary independent and identically distributed (i.i.d.) channel system, framing it as AoI bandits with AoI regret as the optimization metric. Qian et al. [18] established policy-independent lower bounds on the average AoI for multiple users and channels. These studies assumed that the channel state follows a stationary but unknown distribution, or that the channel undergo smooth fading.
@@
However, practical wireless environments exhibit inherently non-stationary scenario[19], manifested through phenomena such as path fading, user mobility in telecommunication networks [20], and hostile jamming in extremely non-stationary communication scenarios [21]. In non-stationary channels, the channel statistical distribution may change abruptly[22]. In federated learning systems, these non-stationary channels led to unpredictable client participation failures during global aggregation, which in turn increases the AoI of client updates. Mandal et al. [21] studied the multi-users scheduling problem and proposed a greedy scheduling policy to minimize AoI of users in adversarial non-stationary environments. For piecewise-stationary channels, the Combinatorial Upper Confidence Bound (CUCB) algorithm [23] was proposed to solve the combinatorial multi-armed bandit problem by balancing exploration and exploitation of unknown channels. Zhou et al. [24] proposed a change-point detector for detecting the moment when the channels undergo sudden changes.
@@
Non-stationary channels can induce participation imbalance in aggregation, challenging federated learning fairness. This phenomenon was particularly prevalent in non-i.i.d. data scenarios, where the global model exhibited bias towards frequently participating clients with superior channel conditions and transmission success rates. Zhu et al. [25] proposed variance reduction to correct the inequality of participation rates of different clients. MAB-based client selection was used in [26] to balance latency and generalization performance. Federated learning fairness covers both collaborative aspects and performance considerations[27][28]. Collaborative fairness addresses contribution evaluation and reward allocation, considering the impact of non-i.i.d. data distribution on model convergence. Performance fairness targets model prediction bias reduction for enhanced cross-client generalization. Huang et al. [29] introduced a long-term fairness constraint for client selection to ensure that the average participation rate of each client did not fall below the expected guaranteed rate. Lyu et al. [30] introduced a fairness evaluation framework that quantifies client reputation based on validation performance. The study in [31] explored the discrepancy between local model updates and the global model, proposing this difference as an indicator of client reputation. To further refine fairness assessment, work in [32] employed the cosine gradient Shapley value to approximate clients’ marginal utility by analyzing the deviation between local and global updates. Jiang et al. [33] focused on estimating the client’s contribution using both gradient and data space to measure fairness in federated learning systems.
@@
To the best of our knowledge, no research has explored the staleness and fairness problem of federated learning under the conditions where CSI is unavailable and the channels are non-stationary. We presents a MAB-based channel scheduling framework for federated learning in non-stationary wireless environments. The main contributions are summarized as follows:
@@
  • For asynchronous federated learning in non-stationary scenarios, we first conduct theoretical analysis, utilizing AoI to quantify client staleness. Our work provides quantitative analysis on the impact of clients’ AoI and successful participation rates and unifies these two factors to derive our optimization objective for channel scheduling. The channel scheduling problem under non-stationary channel conditions is analyzed within the MAB framework.
@@
  • Our analysis encompasses two non-stationary channel scenarios: the piecewise-stationary case and the extremely non-stationary case. Building on this, we propose channel scheduling strategies based on GLR-CUCB and M-exp3 for these scenarios. Furthermore, we derive the achievable theoretical lower bounds on the AoI regret for both cases and establish the corresponding upper bounds, demonstrating their sublinear growth over time.
@@
The organization of this paper proceeds as follows: Section II introduces the asynchronous federated learning model under non-stationary channels. Section III presents the convergence analysis. Section IV establishes the AoI regret lower bounds for both extremely non-stationary and piecewise-stationary channels, and derives the upper bound under the proposed strategy. Section V details the adaptive channel matching algorithm. Section VI evaluates the experimental results. Section VII concludes our research.
@@ II SYSTEM MODEL
This section presents the asynchronous federated learning procedure and wireless channel model, focusing on two non-stationary channel types: extremely non-stationary and piecewise-stationary channels.
@@ II-A Procedure of Federated Learning
We consider a non-stationary wireless network consisting of $M$ clients, indexed by the set =$\mathcal{M}=\{1,2,\dots,M\}$, and a central server. Each client $i\in\mathcal{M}$ holds a local dataset $\mathcal{D}_{i}$, with $|\mathcal{D}_{i}|$ representing the number of data samples. The clients collaborate with the central to complete privacy-preserving decentralized training. We denote the sample-wise loss function for client $i$ as $\mathcal{F}_{i}(\boldsymbol{w}_{i},x)$, where $\boldsymbol{w}_{i}$ represents the local model parameter of client $i$, and $x\in\mathcal{D}_{i}$ is a data sample from the dataset $\mathcal{D}_{i}$. Then the local loss function for client $i$ is given by:
@@
Due to the unpredictability and non-stationarity of CSI, only a subset of clients may successfully transmit their gradient updates to the central server in each round. We define $S_{t}$ as the set of clients that successfully transmit their updates in round $t$. Furthermore, we introduce AoI to quantify the degree to which clients are out-of-sync in the global aggregation process. Specifically, Let $a_{i}(t)$ denote the AoI of client $i$ in aggregation round $t$ and $h_{i}(t)$ denotes the most recent round in which client $i$ successfully participated in the global aggregation before round $t$. We have
@@
$\bullet$ (Step 1 Global Model Broadcast) To mitigate computational resource wastage and optimize the utilization of local training outcomes, clients that failed to transmit their updates in round $t-1$ retain their gradient updates and do not participate in the new updates in round $t$. Therefore, in round $t$, the server only sends the latest global model parameters $\boldsymbol{w}_{t}$ to the clients in $S_{t-1}$. Given the sufficient transmission capacity of the central server, we assume that all transmissions over the downlink channel are always successful.
@@
Clients that failed to participate in the previous round of global aggregation will pause local updates until they successfully transmit their latest parameters to the central server. The central server assigns each client a channel to transmit its model update $\tilde{\boldsymbol{G}}_{i,t}$ to the server.
$\bullet$ (Step 4 Global Model Update) Only the model updates from clients with favorable channel conditions are successfully transmitted to the server. The global model can thus be updated as follows:
@@
where $\zeta_{i}^{t}$ denotes the aggregation weight of client $i$ in round $t$. We rewrite Equation (4) and consequently the AoI of each client $i,\forall i\in\mathcal{M}$ is updated as
@@
We assumpt that $a_{i}(0)=1$ for all clients $i\in\{1,2,\cdots,M\}$ in round $=0$. The above steps are repeated until round $T$ is reached.
@@ II-B Non-stationary Wireless network
In this wireless network, we equally divide the spectrum into $N$ orthogonal sub-channels ($N\geq M$), indexed by the set $\mathcal{N}=\{1,2,\dots,N\}$. Each client periodically attempts to transmit the time-sensitive local updates to the central server over one of the $N$ channels.
The central server schedules distinct sub-channels to clients based on a scheduling policy, preventing collisions within each round. Sub-channels are modeled as Bernoulli channels with state Good (1) or Bad (0) at any round $t$. Clients cannot obtain real-time CSI or prior knowledge of the channels’ statistical properties.
In the stationary setting, channel states are independent between channels over $T$ rounds. Let $\mu_{k}$ denote the mean state of channel $k\in\{1,2,\cdots,N\}$, which is unknown to client $i\in\{1,2,\cdots,M\}$ and constant throughout $T$. Without loss of generality, we assume $\mu_{1}>\mu_{2}>\cdots>\mu_{N}$, where $\{1,\cdots,M\}$ are the $M$-best channels and $\{M+1,\cdots,N\}$ are the $N-M$-worst channels.
In the non-stationary setting, channel states may fluctuate due to factors like malicious attacks or collisions [21]. In the extremely non-stationary scenario, an adversary pre-determines the state sequence of each channel (Good or Bad) without relying on any statistical assumptions[34]. For piecewise-stationary channels, we assume the central server has some prior knowledge, where the mean channel state $\mu_{k}$ for each $k=1,2,\dots,N$ remains constant over certain intervals and changes only at unknown rounds.